From the given table calculate:

  1. Tax on dividend at 40%

Firm A

Firm B

Next year’s price

$140.50

$140.50

Dividend

$ 12

$ 13.00

Total pretax payoff

$140.50

$140.50

Today’s stock price

$300

$ 95.78

Capital gain

$ 13.50

$ 7.78
Answer #1

Tax rate is 40%

Tax on Dividend for Firm A=12*40%=4.80
Tax on Dividend for Firm B=13*40%=5.20